### v2

# Curved representational Bregman divergences and their applications

Abstract
By analogy to curved exponential families in statistics, we define curved Bregman divergences as Bregman divergences restricted to nonlinear parameter subspaces. We show that the barycenter of a finite weighted set of parameters under a curved Bregman divergence amounts to the right Bregman projection onto the nonlinear subspace of the barycenter with respect to the full Bregman divergence. We demonstrate the significance of curved Bregman divergences with two examples: (1) symmetrized Bregman divergences and (2) the Kullback-Leibler divergence between circular complex normal distributions. We then consider monotonic embeddings to define representational curved Bregman divergences and show that the $\alpha$-divergences are representational curved Bregman divergences with respect to $\alpha$-embeddings of the probability simplex into the positive measure cone. As an application, we report an efficient method to calculate the intersection of a finite set of $\alpha$-divergence spheres.
Keywords: Bregman divergences curved exponential family monotonic embeddings $\alpha$-divergences centroids Bregman projection space of spheres.
1 Introduction and contributions
[…]
$B_{F}(\theta:\theta^{\prime}):=F(\theta)-F(\theta^{\prime})-{\langle\theta-\theta^{\prime},\nabla F(\theta^{\prime})\rangle},$ (1)
where $\theta\in\Theta$ and $\theta^{\prime}\in\mathrm{int}(\Theta)$ (the topological interior of $\Theta$). Bregman divergences are non-negative dissimilarity measures (i.e., $B_{F}(\theta:\theta^{\prime})\geq 0$ with equality if and only if $\theta=\theta^{\prime}$) which generalize both the squared Mahalanobis distance and the Kullback-Leibler divergence. Their axiomatization was studied by Csiszár [11].
We consider a class of well-behaved Bregman generators which are of Legendre type [24, 25]: A function $F:\Theta\rightarrow\mathbb{R}$ is of Legendre-type if (i) $\Theta$ is topologically open and (ii) $\lim_{\theta\rightarrow\partial\Theta}\|\nabla F(\theta)\|=\infty$ where $\partial\Theta$ denotes the topological boundary of domain $\Theta$.
When the Bregman generator is of Legendre type [24], the following holds: (a) there is a one-to-one correspondence between the primal parameter $\theta$ and the corresponding dual parameter $\eta$ defined by $\eta(\theta)=\theta^{*}=\nabla F(\theta)$, and (b) the Legendre-Fenchel convex conjugate yields a Legendre-type function: $(H,F^{*})$ given by $F^{*}(\eta)={\langle\eta,\nabla F^{-1}(\eta)\rangle}-F(\nabla F^{-1}(\eta))$, where the open gradient domain is denoted by $H=\{\nabla F(\theta){\ :\ }\theta\in\Theta\}$, and (c) the Legendre-Fenchel transform is an involution, i.e., ${(\Theta,F)^{*}}^{*}=(H,F^{*})^{*}=(F,\Theta)$. Legendre-type Bregman divergences satisfy the biduality identity [29] $B_{F}(\theta_{1}:\theta_{2})=B_{F^{*}}^{\star}(\eta_{1}:\eta_{2})$ where $D^{\star}(x:y)=D(y:x)$ is the reference duality.
The paper is organized as follows: In §2, we define curved Bregman divergences (Definition 2). A characterization of the curved Bregman centroid by Bregman projection is given in §2.2 (Theorem 1). Two examples illustrating curved Bregman divergences are described: First, symmetrized Bregman divergences in §2.3 and second, the Kullback-Leibler geometry of complex circular normal distributions in §2.4. In §3, we further introduce a representation function to define representational curved Bregman divergence (Definition 3). We prove that $\alpha$-divergences are representational curved Bregman divergences (Proposition 1). Finally, as an application, we show how to compute the intersection of a finite set of $\alpha$-divergence spheres in §4 (Proposition 3).
Figure 1: Curved exponential family.
2 Curved Bregman divergences
2.1 Definition
Let $(\mathcal{X},\mathcal{A},\mu)$ be a measure space with $\mathcal{X}$ the sample space, $\mathcal{A}$ the $\sigma$-algebra, and $\mu$ a positive measure. Let us denote by $\mathcal{P}_{\mu}$ the set of probability measures dominated by $\mu$ admitting Radon-Nikodym densities. The notion of curved exponential families in statistics [13, 1] is defined as follows:
Definition 1 (Curved exponential family, Figure 1)
Let $U$ and $V$ be two finite-dimensional vector spaces. Consider a full exponential family $\mathcal{P}_{\Theta}=\{p_{\theta}=\frac{P_{\theta}}{\mathrm{d}\mu}{\ :\ }\theta\in\Theta\subset V\}\subset\mathcal{P}_{\mu}$ with $\dim(\Theta)=m$. Then any injective mapping $u:\mathcal{U}\subset U\rightarrow\Theta,u\mapsto\theta(v)$ with $\dim(\mathcal{U})<\dim(\Theta)$ defines a sub-dimensional statistical model $\mathcal{P}_{\Theta_{|\mathcal{U}}}=\{P_{\theta}(u){\ :\ }u\in\mathcal{U}\}$ that is called a curved exponential family.
Example 1 ([28])
Consider a domain $\mathcal{U}$ of dimension $m^{\prime}$ and an injective linear mapping $\Sigma(u)$ from $\mathcal{U}$ to $\mathrm{Sym}^{++}(d,\mathbb{R})$, the set of $d\times d$ symmetric positive-definite matrices. Then the statistical model $\mathcal{P}^{\prime}=\left\{p_{\mu,\Sigma(u)}{\ :\ }\mu\in\mathbb{R}^{d},u\in\mathcal{U}\right\}$ is a curved exponential family [28] with natural parameter $\theta=\Sigma(u)\mu$ for $u\in\mathcal{U}$. For example, one can choose $\mathcal{U}=\mathbb{R}_{>0}$ (positive reals) with $\Sigma(u)=u\,I$ ($m^{\prime}=1$), where $I$ denotes the $d\times d$ identity matrix.
The notion of curved exponential families allows to consider embedded statistical submanifolds of arbitrary topologies (e.g., Fisher’s normal circle model [14]).
By analogy, let us define the notion of curved Bregman divergences:
Definition 2 (Curved Bregman divergence)
Let $F:\Theta\rightarrow\mathbb{R}$ be a strictly convex and differentiable function defined on the domain $\Theta=\mathrm{dom}(F)\subset\mathbb{R}^{m}$ and $\mathcal{U}\subset\Theta$ with $m^{\prime}=\dim(\mathcal{U})<m$ and $\mathcal{U}=\{\theta(u){\ :\ }u\}$. Then the curved Bregman divergence induced by $F$ on $\mathcal{U}$ is defined by: $\forall u_{1}\in\mathcal{U},u_{2}\in\mathrm{relint}(\mathcal{U})$,
$B_{F}(u_{1}:u_{2}):=F(\theta(u_{1}))-F(\theta(u_{2}))-{\langle\theta(u_{1})-\theta(u_{2}),\nabla_{u}(F(\theta(u_{2}))\rangle},$ (2)
where $\mathrm{relint}(\mathcal{U})$ is the topologically relative interior [6] of $\mathcal{U}$.
Figure 2: The curved Bregman centroid $\theta(\bar{u})$ amounts to the right Bregman projection of the unconstrained Bregman centroid $\bar{\theta}$ onto the subspace $\mathcal{U}$.
2.2 Curved Bregman centroids
We show that the right curved Bregman centroid is characterized by the Bregman projection [16] of the Bregman centroid onto $\mathcal{U}$:
Theorem 1 (Curved Bregman centroid/barycenter)
Let $\theta_{i}=\theta(u_{i})$’s be $n$ weighted parameters of $\mathcal{U}$ with weight vector $w\in\Delta_{n-1}$ (the $(n-1)$-dimensional standard simplex). Then the barycenter in $\mathcal{U}$ with respect to the curved Bregman divergence amounts to the Bregman projection of the center of mass $\bar{\theta}=\sum_{i}w_{i}\theta_{i}$ (right Bregman barycenter) onto $\mathcal{U}$:
$\arg\min_{u\in\mathcal{U}}\sum_{i=1}^{n}w_{i}\,B_{F}(\theta_{i}:\theta(u))=\arg\min_{u\in\mathcal{U}}\,B_{F}(\bar{\theta}:\theta(u)).$ (3)
Proof:
$\min_{u\in\mathcal{U}}\sum_{i=1}^{n}w_{i}\,B_{F}(\theta_{i}:\theta(u))$ $=$ $\sum_{i=1}^{n}w_{i}\left(F(\theta_{i})-F(\theta(u))-{\langle\theta_{i}-\theta(u),\nabla F(\theta(u))\rangle}\right),$
$\equiv$ $-F(\theta(u))-{\langle\bar{\theta}-\theta(u),\nabla F(\theta(u))\rangle},$
$\equiv$ $F(\bar{\theta})-F(\theta(u))-{\langle\bar{\theta}-\theta(u),\nabla F(\theta(u))\rangle},$
$=$ $B_{F}(\bar{\theta}:\theta(u)).$
$\square$
Figure 2 illustrates the right Bregman projection with orthogonality defined with respect to the Hessian metric [26, 27] $\nabla^{2}F(\theta)$ to find the curved Bregman centroid. Theorem 1 generalizes the result of Amari and Nagaoka [4] of the maximum likelihood estimators on curved exponential families. Notice that although the full Bregman centroid $\bar{\theta}$ is unique, the curved Bregman centroid(s) may not be unique when the projection is not unique: e.g., consider $\theta_{i}$ on a unit circle with $\bar{\theta}$ coinciding with the circle center (see also Fisher’s normal circle model [14]). Then the projection can be any point of the circle. That is, let $B_{F}(\theta_{1}:\theta_{2})=\frac{1}{2}{\langle\theta_{2}-\theta_{1},\theta_{2}-\theta_{1}\rangle}=\frac{1}{2}\|\theta_{1}-\theta_{2}\|^{2}$ be the Bregman divergence obtained for $F(\theta)=\frac{1}{2}{\langle\theta,\theta\rangle}$. Consider $U=\mathbb{R}$ and $\mathcal{U}=[0,2\pi)\subset U$ with $\theta(u)=(\cos\theta,\sin\theta)\in\mathbb{R}^{2}=:V$. We have $B_{F}(u_{1}:u_{2})=\frac{1}{2}((\cos u_{2}-\cos u_{1})^{2}+(\sin u_{2}-\sin u_{1})^{2})$. Let $u_{1}=0$ and $u_{2}=\pi$ with $\theta_{1}=(1,0)$ and $\theta_{2}=(-1,0)$. We have $\bar{\theta}=(0,0)$ and $\arg\min_{u}B_{F}((0,0):\theta(u))=\arg\min_{u}\frac{1}{2}\|\cos^{2}\theta+\sin^{2}\theta\|^{2}=\arg\min_{u}\frac{1}{2}=\mathcal{U}$, i.e., the curved Bregman centroid is the full circle $\{\theta(u){\ :\ }u\in\mathcal{U}\}$.
Corollary 1 ([4])
Let $x_{1},\ldots,x_{n}$ be $n$ i.i.d. observations of a curved exponential family $\{p_{\theta(u)}(x){\ :\ }u\in\Theta_{\mathcal{U}}\}$ with sufficient statistic vector $t(x)$ and cumulant function $F(\theta)$. Then the maximum likelihood estimator amounts to a curved Bregman centroid on $t(x_{1}),\ldots,t(x_{n})$ with respect to the convex conjugate $F^{*}(\eta)$. That is, the right dual Bregman projection of the observed point $\bar{t}=\frac{1}{n}\sum_{i=1}^{n}t(x_{i})$ onto the moment submanifold $\{\eta(u)=\nabla F(\theta(u))\}$:
$\eta(\hat{u})=\min_{u\in\mathcal{U}}B_{F^{*}}(\bar{t}:\eta(u)).$
Proof: For regular full exponential families (i.e., $\theta(u)$ the identity function), using the duality between regular exponential families and regular Bregman divergences [6], we have [15]:
$\max_{\theta\in\Theta}\prod_{i=1}^{n}p_{\theta}(x_{i})\equiv\min_{\eta}\sum_{i=1}^{n}B_{F^{*}}(t(x_{i}):\eta).$
Let $\eta(u)=\nabla F(\theta(u))$. It follows that we have
$\max_{u}\prod_{i=1}^{n}p_{\theta(u)}(x_{i})\equiv\min_{u}B_{F^{*}}(\bar{t}:\eta(u)),$
where $\bar{t}=\frac{1}{n}\sum_{i=1}^{n}t(x_{i})$ is the observed point [4]. $\square$
Let us report now two examples of curved Bregman divergences.
2.3 Symmetrized Bregman divergences as curved Bregman divergences
Consider the symmetrized Bregman divergence:
$S_{F}(\theta_{1},\theta_{2})$ ::=:= $B_{F}(\theta_{1}:\theta_{2})+B_{F}(\theta_{2}:\theta_{1}),$
$=$ $B_{F}(\theta_{1}:\theta_{2})+B_{F^{*}}(\nabla F(\theta_{1}):\nabla F(\theta_{2})),$
$=$ $B_{F_{\xi}}(\xi(\theta_{1}):\xi(\theta_{2})),$
where $\theta_{i}\in\Theta\subset\mathbb{R}^{m}$, $\xi(\theta):=(\theta,\nabla F(\theta))\in\mathbb{R}^{2m}$ and
$F_{\xi}(\theta):=F(\theta)+F^{*}(\nabla F(\theta)),$
where $F^{*}(\eta)={\langle\theta,\eta\rangle}-F(\theta)$ denotes the convex conjugate of $F(\theta)$. The symmetrized Bregman divergence $S_{F}$ (also called Jeffreys-Bregman divergence [17]) is a curved Bregman divergence $B_{F_{\xi}}$ with respect to the generator $F_{\xi}$ and the (usually non-linear) subspace $\mathcal{U}=\{u(\theta)=(\theta,\nabla F(\theta)){\ :\ }\theta\in\Theta\}$. Thus the symmetrized Bregman centroid [20] of a weighted parameter set $\{\theta_{i}\}$ is the projection of the centroid $\bar{\xi}=(\sum w_{i}\theta_{i},\sum w_{i}\nabla F(\theta_{i}))$ onto $\mathcal{U}=\{(\theta,\nabla F(\theta)){\ :\ }\theta\in\Theta\}$:
$\arg\min_{\theta\in\Theta}\sum_{i}w_{i}\,S_{F}(\theta_{i},\theta)$ $=$ $\arg\min_{\theta\in\Theta}\sum_{i}w_{i}\,B_{F_{\xi}}(\xi_{i}:\xi(\theta)),$
$=$ $\arg\min_{u\in\mathcal{U}}B_{F_{\xi}}(\bar{\xi}:u),$
where $\xi_{i}=(\theta_{i},\eta_{i}=\nabla F(\theta_{i}))$ and $\bar{\xi}=\sum_{i}w_{i}\xi_{i}$.
2.4 Kullback-Leibler divergence between circular complex normal distributions
The $d$-variate circular complex normal distribution [23] $\mathcal{CN}_{d}(\mu_{\mathbb{C}},S_{\mathbb{C}})$ can be handled as a $2d$-real normal distribution $N_{2d}([\mu_{\mathbb{C}}]_{\mathbb{R}},\frac{1}{2}[S_{\mathbb{C}}]_{\mathbb{R}})$ where $[z=a+ib]_{\mathbb{C}}=(a,b)$ and $[M=A+iB]_{\mathbb{C}}={\left[\begin{array}[]{ll}A&-B\cr B&A\end{array}\right]}$, where $a,b\in\mathbb{R}$ and $A,B\in\mathbb{R}^{d\times d}$. Let $N(m_{\mathbb{C}},S_{\mathbb{C}})$ and $N(m_{\mathbb{C}}^{\prime},S_{\mathbb{C}}^{\prime})$ be two circular complex normal distributions. The Kullback-Leibler divergence between those two distributions amounts to a KLD between their corresponding real normal distributions $N_{2d}(\mu,\Sigma)$ and $N_{2d}(\mu^{\prime},\Sigma^{\prime})$ where $\mu=[m_{\mathbb{C}}]_{\mathbb{C}}$, $\Sigma=[S_{\mathbb{C}}]_{\mathbb{C}}$, and $\mu^{\prime}=[m^{\prime}_{\mathbb{C}}]_{\mathbb{C}}$, $\Sigma^{\prime}=[S_{\mathbb{C}}^{\prime}]_{\mathbb{C}}$. Since the KLD between two densities of an exponential family amounts to a reverse Bregman divergence between their natural parameters [5], we have
$D_{\mathrm{KL}}[p_{m_{\mathbb{C}},S_{\mathbb{C}}}:p_{m_{\mathbb{C}}^{\prime},S_{\mathbb{C}}^{\prime}}]=D_{\mathrm{KL}}[p_{\mu,\Sigma}:p_{\mu^{\prime},\Sigma^{\prime}}]=B_{F}(\theta^{\prime}:\theta).$ (4)
The Gaussian family $\{N(\mu,\Sigma)\}$ is an exponential family for natural parameter $\theta=(\Sigma^{-1}\mu,\frac{1}{2}\Sigma^{-1})$ and cumulant function (or log-partition function) $F(\theta)=F(\theta_{1},\theta_{2})=\frac{1}{2}\left(d\log\pi-\log\det(\theta_{2})+\frac{1}{2}\theta_{1}^{\top}\theta_{2}^{-1}\theta_{1}\right)$. Thus the family of $d$-variate circular complex normal distributions is a curved exponential family of $2d$-variate real normal distributions with
$\mathcal{U}=\left\{(v,M){\ :\ }v\in\mathbb{R}^{2d},M={\left[\begin{array}[]{ll}A&-B\cr B&A\end{array}\right]},A\in\mathbb{R}^{d\times d}\succ 0,B\in\mathbb{R}^{d\times d}\succ 0\right\}.$
When $\Theta_{\mathcal{U}}=\{u(\theta){\ :\ }\theta\in\Theta\}$ is the restriction of a natural parameter space to an affine subspace, the curved Bregman divergence is a sub-dimensional Bregman divergence.
Example 2
The Kullback-Leibler divergence (KLD) $D_{\mathrm{KL}}^{+}$ extended to positive arrays [11] $\Theta=\mathbb{R}_{>0}^{m}$ (positive orthant cone) is a separable Bregman divergence obtained for the generator $F_{\mathrm{KL}^{+}}(u)=\sum_{i=1}^{m}u_{i}\log u_{i}$: $D_{\mathrm{KL}}^{+}(\theta:\theta^{\prime})=\sum_{i=1}^{m}\theta_{i}\log\frac{\theta_{i}}{\theta^{\prime}_{i}}+\theta_{i}^{\prime}-\theta_{i}$. Consider the hyperplane $H_{\Delta}:\sum_{i=1}^{m}x_{i}=1$. Then $\Theta_{|H_{\Delta}}=\Delta_{m}$ is the standard simplex, and $u(x)=(x_{1},\ldots,x_{m-1})\in\Delta_{m-1}$ is the moment parameter of the simplex exponential family. Then the curved Bregman divergence $B_{F_{\mathrm{KL}^{+}}}(u(\theta):u(\theta^{\prime}))$ which corresponds to the KLD defined on the standard simplex amounts to a sub-dimensional Bregman divergence: $B_{F_{\mathrm{KL}^{+}}}(u(\theta):u(\theta^{\prime}))=B_{F_{\mathrm{KL}}}(\alpha:\alpha^{\prime})$ with $\alpha=u(\theta)$, $\alpha^{\prime}=u(\theta^{\prime})$ in $\Delta_{m}$ for the non-separable generator $F_{\mathrm{KL}}(\alpha)=\sum_{i=1}^{m-1}\alpha_{i}\log\alpha_{i}+(1-\sum_{i=1}^{m-1}\alpha_{i})\log(1-\sum_{i=1}^{m-1}\alpha_{i})$.
3 Curved representational Bregman divergences
3.1 Definition
Definition 3
Let $B_{F}$ be a Bregman divergence for the Legendre-type generator $F:\Theta\subset V\rightarrow\mathbb{R}$. Consider a diffeomorphism $R:\Theta\rightarrow\Theta,\theta\mapsto r=R(\theta)$ called the representation function. Then $B_{F}(R(\theta_{1}):R(\theta_{2}))=B_{F}(r_{1}:r_{2})$ is termed a representational Bregman divergence [21].
Note that when $R$ is the identity function (i.e., $R(\theta)=\theta$), representational Bregman divergences become ordinary Bregman divergences. Examples of representational Bregman divergences are met when dealing with the Kullback-Leibler divergence (KLD) between two densities of a generic exponential family [7], not necessarily a natural exponential family [5, 18]. Let $(\mathcal{X},\mathcal{F},\mu)$ be a measure space and =$\mathcal{E}=\{p_{\lambda}(x)=\exp({\langle t(x),\theta\rangle}(\lambda))-F(\theta(\lambda))\mathrm{d}\mu\}$ an exponential family with sufficient statistic vector $t(x)$, natural parameter $\theta(\lambda)$ and cumulant/log-normalizer/log-partition function $F(\theta)$. It is well-known that the KLD $D_{\mathrm{KL}}(p_{\lambda_{1}}:p_{\lambda_{2}})=\int p_{\lambda_{1}}(x)\log\frac{p_{\lambda_{1}}(x)}{p_{\lambda_{2}}(x)}\mathrm{d}\mu$ between two densities $p_{\lambda_{1}}$ and $p_{\lambda_{2}}$ of an exponential family amounts to a reverse representational Bregman divergence:
$D_{\mathrm{KL}}(p_{\lambda_{1}}:p_{\lambda_{2}})=B_{F}(\theta(\lambda_{2}):\theta(\lambda_{1})).$
Here, the natural parameterization $\theta(\lambda)$ plays the role of the representation function.
Proposition 1
The $\alpha$-divergences [3] $D_{\alpha}^{+}(q_{1}:q_{2})$ extended to the positive measures $q_{1}$ and $q_{2}$ of $\mathbb{R}_{>0}^{m}$:
$D_{\alpha}^{+}(q_{1}:q_{2})=\left\{\begin{array}[]{ll}\frac{4}{1-\alpha^{2}}\,\sum_{i=1}^{m}\left(\frac{1-\alpha}{2}q_{1}+\frac{1+\alpha}{2}q_{2}-q_{1}^{\frac{1-\alpha}{2}}\,q_{2}^{\frac{1+\alpha}{2}}\right),\alpha\in\mathbb{R}\backslash\{-1,1\}\\ {D_{\mathrm{KL}}^{\star}}^{+}(q_{1}:q_{2})=D_{\mathrm{KL}}^{+}(q_{2}:q_{1})=\sum_{i=1}^{m}q_{2}^{i}\log\frac{q_{2}^{i}}{q_{1}^{i}}+q_{1}^{i}-q_{2}^{i}&\alpha=1\\ D_{\mathrm{KL}}^{+}(q_{1}:q_{2})=\sum_{i=1}^{m}q_{1}^{i}\log\frac{q_{1}^{i}}{q_{2}^{i}}+q_{2}^{i}-q_{1}^{i}&\alpha=-1.\end{array}\right.$
is a representational Bregman divergence:
$D_{\alpha}^{+}(q_{1}:q_{2})=B_{F_{\alpha}}(R_{\alpha}(q_{1}):R_{\alpha}(q_{2}))=B_{F_{-\alpha}}(R_{-\alpha}(q_{2}):R_{-\alpha}(q_{1})),$ (5)
where $F_{\alpha}(r)=\sum_{i=1}^{m}\,f_{\alpha}(r_{i})$ is the separable Bregman generator on the representations induced by the potential function
$f_{\alpha}(x)=\left\{\begin{array}[]{ll}\frac{2}{1+\alpha}\,\left(\frac{1-\alpha}{2}x\right)^{\frac{2}{1-\alpha}},&\alpha\not=1\\ x\exp(x),&\alpha=1.\end{array}\right.$
and $r=R_{\alpha}(q)=(r_{\alpha}(q_{1}),\ldots,r_{\alpha}(q_{m}))$ is the representation function induced by
$r_{\alpha}(x)=\left\{\begin{array}[]{ll}\frac{2}{1-\alpha}\left(x^{\frac{1-\alpha}{2}}-1\right),&\alpha\not=1,\\ \log x,&\alpha=1.\end{array}\right.$
Furthermore, $D_{\alpha}^{+}$ can be expressed as an equivalent representational Fenchel-Young divergence:
$D_{\alpha}^{+}(q_{1}:q_{2})=Y_{F_{\alpha},F_{-\alpha}}(R_{\alpha}(q_{1}):R_{-\alpha}(q_{2})),$ (6)
where $F_{-\alpha}$ is the convex conjugate $F_{\alpha}^{*}$ of $F_{\alpha}$ and $R_{-\alpha}$ is the dual representation [21].
Amari [2] proved that the intersection of the class of $f$-divergences extended on positive measures with the class of Bregman divergences are exactly the extended $\alpha$-divergences.
4 Application: Intersection of $\alpha$-divergence spheres
It is well-known that the intersection of two $m$-dimensional Euclidean spheres $\Sigma_{1}$ and $\Sigma_{2}$ is yet another $(m-1)$-dimensional Euclidean sphere. This section first deals with the intersection of two Bregman spheres induced by the same Bregman generator yielding another type of $(d-1)$-dimensional Bregman sphere (i.e., a sub-dimensional Bregman divergence sphere).
Let $\sigma_{F}(\theta,r)$ and $\sigma^{\star}_{F}(\theta,r)$ be the right and left Bregman spheres, respectively:
$\sigma_{F}(\theta,r):=\left\{\theta^{\prime}\in\Theta{\ :\ }B_{F}(\theta^{\prime}:\theta)=r\right\},\quad\sigma_{F}^{\star}(\theta,r):=\left\{\theta^{\prime}\in\Theta{\ :\ }B_{F}(\theta:\theta^{\prime})=r\right\},$
[…]
$\sigma_{F}^{\star}(\theta,r)=\left(\sigma_{F^{*}}(\eta=\theta^{*},r)\right)^{*},\quad\sigma_{F^{*}}^{\star}(\eta,r)=\left(\sigma_{F}(\eta^{*},r)\right)^{*}.$
We now explain the space of Bregman spheres [8]: For a parameter $\theta$, let $\hat{\theta}:=(\theta,F(\theta))$ be the vertical lifting of $\theta$ onto the epigraph :=$\mathcal{F}:=\{(\theta,F(\theta){\ :\ }\theta\in\Theta)\}$ and $\downarrow\hat{\theta}=\theta$ be the vertical projection so that $\downarrow(\hat{\theta})=\theta$. Let $\Sigma=\sigma_{F}(\theta,r)$ be a right Bregman sphere. The pointwise lifted right Bregman sphere $\hat{\Sigma}=\hat{\sigma}_{F}(\theta,r)$ is supported by a non-vertical hyperplane $H_{\Sigma}$ of equation [8]:
$H_{\Sigma}=H_{\theta,r}:y={\langle\theta^{\prime}-\theta,\nabla F(\theta)\rangle}+F(\theta)+r.$
That is, we have $\Sigma=\downarrow(H_{\Sigma}\cap\mathcal{F})$.
Conversely, the intersection of any non-vertical hyperplane $H:y={\langle\theta_{a},\theta\rangle}+b$ with $\mathcal{F}$ projects vertically on $\Theta$ as a Bregman sphere with center $\theta_{H}=\nabla F^{-1}(\theta_{a})$ and radius $r_{H}={\langle\theta_{a},\theta_{H}\rangle}-F(\theta_{H})+b$.
The supporting hyperplane of the right Bregman sphere (Figure 3, blue) is obtained by lifting the center $\theta$ onto the epigraph, taking the tangent plane at $(\theta,F(\theta))$ and translating that tangent plane vertically by $r\geq 0$. The left Bregman sphere (Figure 3, red) is obtained by first illuminating [12] from $(\theta,F(\theta)-r)$ the epigraph $\mathcal{F}$, and then projecting vertically the illuminated portion of $\mathcal{F}$ onto $\Theta$. Let $\mathcal{F}^{\circ}$ denote the open interior of the epigraph $\mathcal{F}$. Then the illuminated portion of the graph =$\partial\mathcal{F}=\{(\theta,F(\theta)){\ :\ }\theta\in\Theta\}$ is given by $\left({\overline{\mathrm{CH}}}\left(\{(\theta,F(\theta)-r)\}\cup\mathcal{F}\right)\backslash\mathcal{F}^{\circ}\right)\cap\partial\mathcal{F}$, where ${\overline{\mathrm{CH}}}$ denotes the closed convex hull operator. Figure 4 illustrates the left and right Bregman sphere obtained for the scalar Shannon negentropy function $F_{S}(\theta)=\theta\log\theta$.
Figure 3: Construction of Bregman spheres from the Bregman generator epigraph: Right Bregman sphere $\sigma_{F}(\theta,r)$ (blue) and left Bregman sphere $\sigma_{F}^{\star}(\theta^{\prime},r^{\prime})$ (red).
Refer to caption Refer to caption Refer to caption
Figure 4: Illustrations of the lifting of left and right Bregman spheres on the potential function induced by the Shannon neg-entropy function. Right Bregman sphere is obtained as the vertical projection of the hyperplane with the potential function (left figure). Left Bregman sphere is given as the intersection of the graph of the potential function with the convex hull including the point (i.e., the illuminated part of the potential function from that point, middle figure). Superposition of both right and left Bregman spheres (right figure).
The lifting of Bregman spheres $\sigma_{F}$ to the potential epigraph $\mathcal{F}$ allows one to build efficient algorithms for computing the intersection of Bregman spheres:
Proposition 2 ([8])
The intersection of two Bregman spheres $\Sigma_{1}$ and $\Sigma_{2}$ of dimension $m$ is a sub-dimensional Bregman sphere $\Sigma_{12}$ of dimension $m-1$.
Proof: Let $\Sigma_{1}=\sigma_{F}(\theta_{1},r_{1})=\{\theta^{\prime}{\ :\ }B_{F}(\theta^{\prime}:\theta_{1})=r_{1}\}$ and $\Sigma_{2}=\sigma_{F}(\theta_{2},r_{2})=\{\theta^{\prime}{\ :\ }B_{F}(\theta^{\prime}:\theta_{2})=r_{2}\}$ be two right Bregman spheres. We have $\Sigma_{1}=\downarrow(H_{\Sigma_{1}}\cap F)$ and $\Sigma_{2}=\downarrow(H_{\Sigma_{2}}\cap F)$. Thus, we have
$\Sigma_{1}\cap\Sigma_{2}$ $=$ $\downarrow(H_{\Sigma_{1}}\cap F)\cap\downarrow(H_{\Sigma_{1}}\cap F)=\downarrow((H_{\Sigma_{1}}\cap H_{\Sigma_{2}})\cap F).$
Let $H_{12}=H_{\Sigma_{1}}\cap H_{\Sigma_{2}}$ be a flat (affine subspace) of dimension $m-1$, and let $H_{12}^{|}$ be the vertical flat (with respect to $y$-axis) of $\mathbb{R}^{m+1}$ passing through $H_{12}$. Then $\mathcal{F}_{12}=\mathcal{F}\cap H_{12}^{|}$ is the epigraph of a Bregman generator $F_{12}$. Thus the intersection of $\Sigma_{1}$ with $\Sigma_{2}$ is a Bregman sphere with respect to generator $F_{12}$. $\square$
Figure 5 illustrates the representation of 2D Euclidean spheres by 3D planes cutting the paraboloid of revolution. The boolean algebra of the union and intersection of Bregman balls can be manipulated as intersections of halfspaces cutting the epigraph of the potential function [8].
Refer to caption Refer to caption
Figure 5: Representing Bregman spheres $\Sigma$ as intersections of hyperplanes $H_{\Sigma}$ with a potential function $\mathcal{F}$ (paraboloid of revolution for Euclidean spheres). Left: Two 2D Bregman spheres represented by 3D planes $H_{\Sigma_{1}}$ and $H_{\Sigma_{2}}$. Right: Their union $\Sigma_{1}\cup\Sigma_{2}$ can be visualized by looking at the interior of the paraboloid from the top. Their radical axis is the vertical projection of the intersection $H_{\Sigma_{1}}\cap H_{\Sigma_{2}}$.
Now consider $\alpha$-divergence spheres ($\alpha$-spheres for short): Let $\sigma_{\alpha}(\theta,r):=\left\{\theta^{\prime}\in\Theta{\ :\ }D_{\alpha}(\theta^{\prime}:\theta)=r\right\}$ be the right $\alpha$-sphere of center $\theta$ and radius $r$. Then the left $\alpha$-divergence sphere $\sigma_{\alpha}^{\star}(\theta,r):=\left\{\theta^{\prime}\in\Theta{\ :\ }D_{\alpha}(\theta:\theta^{\prime})=r\right\}=\sigma_{-\alpha}(\theta,r)$. Since $\alpha$-divergences are curved representational Bregman divergences, we may consider the following algorithm to calculate the intersection of $n$ $\alpha$-spheres $\Sigma_{1}=\sigma_{\alpha}(\theta_{1},r_{1}),\ldots,\Sigma_{n}=\sigma_{\alpha}(\theta_{n},r_{n})$: Let $S_{i}=\sigma_{F_{\alpha}}(r_{\alpha}(\theta_{i}),r_{i})$ be the representational Bregman sphere corresponding to $\Sigma_{i}$ for $i\in\{1,\ldots,n\}$, and $\mathcal{F}_{\alpha}$ the epigraph of $F_{\alpha}(\theta)=\frac{2}{1+\alpha}\sum_{i=1}^{m}\left(\frac{1-\alpha}{2}\theta_{i}\right)^{\frac{2}{1-\alpha}}$.
Proposition 3 (Intersection of $\alpha$-spheres)
We have
[…]
The representational Bregman divergence point of view of $\alpha$-divergences (Proposition 1) allows one to extend easily the computational geometry toolbox of Bregman divergences to $\alpha$-divergences. For example, we may consider nearest neighbor query data structures for $\alpha$-divergences following [22], or consider the smallest enclosing $\alpha$-divergence ball following [19]. The circumcenter of the smallest enclosing ball is also called the Chebyshev point [10] and finds application in universal coding in information theory.
5 Summary
In this work, we defined the notion of curved Bregman divergences (Definition 1) by analogy to curved exponential families [1]. We proved that the barycenter of a finite weighted set of points belonging to a nonlinear subspace $\mathcal{U}$ (submanifold) of a Bregman parameter space $\Theta$ amounts to a right Bregman projection of the center of mass onto $\mathcal{U}$ (Theorem 1). We reported several examples of curved Bregman divergences: (1) symmetrized Bregman divergences (Jeffreys-Bregman divergence [17]), (2) Kullback-Leibler divergence between circular complex normal distributions, and (3) Kullback-Leibler divergence between categorical distributions (Example 2). As an application, we handled $\alpha$-divergences as curved representational Bregman divergences and show how to compute the intersection of $\alpha$-spheres (Proposition 1), i.e., spheres defined according to $\alpha$-divergences.